# Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis

### v2

@@ Abstract
CP violation in baryon decays has recently garnered significant attention. Traditionally, CP violation in charmed baryon decays is predicted to be at the order of $O(10^{-4})$. Utilizing recent measurements from BESIII, Belle and Belle II experiments, we conduct a comprehensive global analysis to estimate the contribution of topological diagrams in these processes. Our study reveals that topological diagrams alone cannot fully account for the experimental data. By incorporating new effects that are not included in traditional topological diagrams and exploring a novel way to determine the possible new weak phase, our analysis reveals the potential for significant CP violation in $\Lambda_{c}^{+}\to p\eta/n\pi^{+}$ and $\Xi_{c}^{+}\to\Xi^{0}K^{+}$. This suggests a promising opportunity to observe CP violation for the first time in charmed baryon decays.
@@
In our analysis, we provide a global fit to determine all the IRA amplitudes in Ref. Xing:2023dni . With the help of the equivalence between IRA and TDA, we can not only test the correctness of the topological diagrams but also estimate the value of each diagram in TDA. As a result, we find that although the topological diagrams in TDA are basically consistent with our IRA numerical results, the topological diagram which only includes the $O_{1,2}$ operator cannot perfectly explain our IRA results. Without loss of generality, we can use a new diagram to express the conflicting parts of IRA and TDA and include them in our fit without any assumptions about their source. To analyze the possible CPV introduced by the new diagram, we explored a novel way to determine their weak and strong phases. Our analysis shows that the new diagram may introduce a large CPV and it is promising to be observed in future experiments.
The global analysis of anti-triplet charmed baryon two-body decays with IRA- Using SU(3) flavor symmetry, the anti-triplet charmed baryon, light baryon octet, and light meson octet can be expressed by $3\times 3$ matrices $T_{c\bar{3}}$, $T_{8}$, and $P$, respectively Xing:2023dni . The SU(3) invariant decay amplitudes for anti-triplet charmed baryon two-body decays are constructed as $a_{15}\times\epsilon_{imn}(T_{c\bar{3}})^{[mn]}(H_{\overline{15}})^{\{ik\}}_{j}(\overline{T_{8}})^{j}_{k}P^{l}_{l}$ and $a_{6}\times(T_{c\bar{3}})^{[ik]}(H_{\overline{6}})_{\{ij\}}(\overline{T_{8}})^{j}_{k}P^{l}_{l}$. Here the effective weak interaction Hamiltonian with IRA method is decomposed: $3\otimes\bar{3}\otimes 3=3\oplus 3\oplus\bar{6}\oplus 15$. The nonzero matrix are $H_{\bar{6}}$ and $H_{15}$, which are expressed in Ref. Xing:2023dni . By enumerating all possible matrix combinations, nine independent amplitudes can be constructed as $\{a,b_{6},c_{6},d_{6}\}$ and $\{a^{\prime},b_{15},c_{15},d_{15},e_{15}\}$. Using these nine SU(3) amplitudes, all amplitudes of anti-triplet charmed baryon two-body decays can be expressed.
@@
The assumptions are approximately supported by the numerical results within $1\sigma$ in Table. 1. However, some form factors contradict the assumption, such as $|f_{1,15}|-|f_{2,16}|\sim 0.06$, $|f_{T_{17,18}}|-|f_{19}|\sim 0.02$ and $|g_{17,19}|-|g_{18}|\sim 0.05$. To focus on the largest effects, we ignore the discrepancy involving $f_{19}$ and primarily consider $|f_{1,15}|>|f_{2,16}|$ and $|g_{17,19}|>|g_{18}|$ in the subsequent analysis. This implies that there are other effects in the topology diagrams. For exploring the possible new effects, one defines the form factors $\textbf{f}_{b}$ and $\textbf{g}_{b}$ to absorb these effect as $\textbf{f}_{b}=f_{T_{1}}-f_{T_{2}}$ and $\textbf{g}_{b}=-g_{18}-g_{T_{7}}$. The corresponding topology diagrams $T_{8,9}$ are given in Fig. 1. Then the IRA amplitude is expressed by TDA form factors and $\textbf{f}_{b},\textbf{g}_{b}$ as
@@
$g^{c}_{6}=g_{T_{6}},\quad g^{d}_{6}=3g_{T_{1}}-g_{T_{7}}+\frac{\textbf{g}_{b}}{2},\quad g^{d}_{15}=g_{T_{6}},$ (5)
Since all the diagrams induced by $O_{1,2}$ are included in TDA diagrams $T_{1-7}$, it is reasonable to believe the diagrams $T_{8,9}$ are induced by SU(3) symmetry breaking or other operators like the penguin operators $O_{3-6}$. The inclusion of these additional diagrams suggests the possibility of the involvement of a new weak phase in these processes. In Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis, we find that $g^{c}_{15}=0$, which is consistent with our global fit (Case I) in Table. 1. One can also directly observe that $f^{b}_{6}$ depends only on new effect form factors, which provides us an ideal condition to study whether a different weak phase exists in these processes.
@@
form factors Case II ($\chi^{2}$/d.o.f=0.597)
@@
vector(f) $f^{a}=-0.0368(84)$ $f^{b}_{6}=0.0075(49)$ $f^{c}_{6}=0.017(13)$ $f_{T_{1}}=-0.0119(35)$ $f_{T_{4}}=0.0080(92)$
$f^{d}_{6}=0.006(27)$ $f^{a\prime}=-0.008(332)$ $f_{T_{6}}=-0.018(14)$ $f_{T_{7}}=0.014(13)$
$f^{b}_{15}=0.0201(71)$ $f^{c}_{15}=-0.0055(65)$ $f^{d}_{15}=-0.018(20)$ $\textbf{f}_{b}=-0.0013(58)$
$f^{e}_{15}=-0.0130(61)$
axial-vector(g) $g^{a}=-0.022(38)$ $g^{b}_{6}=0.035(15)$ $g^{c}_{6}=-0.036(21)$ $g_{T_{1}}=-0.024(51)$ $g_{T_{3}}=-0.18(60)$
$g^{d}_{6}=0.03(10)$ $g^{a\prime}=0.13(38)$ $g_{T_{4}}=0.11(12)$ $g_{T_{6}}=0.018(52)$
$g^{b}_{15}=-0.197(24)$ $g^{c}_{15}=-0.024(65)$ $g^{d}_{15}=-0.015(74)$ $g_{T_{7}}=0.02(10)$ weak phase
$g^{e}_{15}=-0.005(55)$ $\phi=-1.16(85)$
@@
According to Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis, the IRA form factor $f^{b}_{6}$ depends solely on the new effect, potentially introducing a new weak phase $\phi$ that could lead to significant CP violation.
In our global analysis, the form factors are defined as
$A^{q}_{6,15}$ $=$ $e^{i\phi_{1}}\bigg{(}\mathcal{R}e(A^{q}_{6,15})+\mathcal{I}m(A^{q}_{6,15})\bigg{)}=|A^{q}_{i}|e^{i\phi_{1}}e^{i\delta^{q}_{i}},$
$f^{b}_{6}$ $=$ $e^{i\phi}\bigg{(}\mathcal{R}e(f^{b}_{6})+\frac{3}{2}\mathcal{I}m(\textbf{f}_{b})\bigg{)}=|f^{b}_{6}|e^{i\phi}e^{i\delta^{b}_{6}},$
$f^{d}_{6}$ $=$ $e^{i\phi_{1}}\bigg{(}\mathcal{R}e(f^{d}_{6})+\mathcal{I}m(f^{d}_{6})+\mathcal{R}e(f^{b}_{6})\bigg{)}$
$-e^{i\phi}\bigg{(}\mathcal{R}e(f^{b}_{6})+\frac{3}{2}\mathcal{I}m(\textbf{f}_{b})\bigg{)},$
$f^{e}_{15}$ $=$ $e^{i\phi_{1}}\bigg{(}\mathcal{R}e(f^{e}_{15})+\mathcal{I}m(f^{e}_{15})+\mathcal{R}e(f^{b}_{6})\bigg{)}$ (6)
$-e^{i\phi}\bigg{(}\mathcal{R}e(f^{b}_{6})+\frac{3}{2}\mathcal{I}m(\textbf{f}_{b})\bigg{)},A=f,g,$
where $q=a,b,c,d,e$. The weak phase $\phi_{1}$ arises from the current-current operator as $\phi_{1}=\arg(V_{cq}^{*}V_{uq^{\prime}})\approx 0,-\pi$.
Although we can obtain the phase of $\textbf{f}_{b}$ in our fit, we cannot distinguish whether the phase is from the weak phase or the strong phase due to the lack of charmed anti-baryon decay data. Fortunately, in the SM, $\Delta c=1$ decay processes inducing CPV at the quark level are categorized into two types, each proportional to the CKM matrix elements $V_{ud}V_{cd}^{*}$ and $V_{us}V_{cs}^{*}$ He:2024pxh . This indicates that the weak phase only enters the Cabibbo-suppressed processes, which provides a new way to determine the strong phase and weak phase of $\textbf{f}_{b}$. Because SU(3) symmetry requires the strong phase in the same form factor corresponding to different processes to be equal Geng:2023pkr ; Zhong:2024qqs , one has the chance to determine the strong phase by the Cabibbo-allowed and doubly Cabibbo-suppressed processes. Based on the strong phase we derived, the weak phase can be further determined by the Cabibbo-suppressed processes.
@@
By introducing the new weak phase $\phi$, one obtains the fit results with $\chi^{2}/d.o.f=0.597$, indicating that our fit is reasonable and the form factors effectively explain the experimental data.
@@
With the determined weak phase $\phi=-1.16\pm 0.85$, one can predict the CP violation ($A_{CP}$) for the Cabibbo-suppressed processes. Given that the QCD penguin operator is likely responsible for this new weak phase, we observe that the fit vaule aligns with expectations for the QCD penguin’s weak phase. Hence, our study implies that the contribution of the QCD penguin may be more substantial than conventionally assumed, necessitating further theoretical investigation in the future.
@@
Using the nontrivial weak phase $\phi$, we predict the values of $A_{CP}$ for the Cabibbo-suppressed channels and found that almost all the predicted $A_{CP}$ values are approximately zero within errors, except for the following processes
$A_{CP}^{\Lambda_{c}^{+}\to p\eta}=-0.047(45),\quad A_{CP}^{\Lambda_{c}^{+}\to n\pi^{+}}=-0.33(28),$
$A_{CP}^{\Xi_{c}^{+}\to\Xi^{0}K^{+}}=-0.39(32).$ (7)
We strongly recommend the experimental measurement of these two processes: $\Lambda_{c}^{+}\to p\eta/n\pi^{+}$ given their measured branching ratios. Additionally, the following processes also have a high probability of being observed in future experiments,
$A_{CP}^{\Xi_{c}^{+}\to\Sigma^{+}\eta}=-0.08(10),\quad A_{CP}^{\Xi_{c}^{+}\to\Lambda\pi^{+}}=-0.23(26),$
$A_{CP}^{\Xi_{c}^{0}\to\Sigma^{0}\pi^{0}}=-0.13(14).$ (8)
@@
By leveraging the equivalence between IRA and TDA methods, our global analysis extends to complex form factors. By determining the strong phase of new effect form factors $\textbf{f}_{b}$, the weak phase of this form factor can be determined in a completely new way with $\phi=-1.16\pm 0.85$. Then we further explore CP violation in anti-triplet charmed baryon two-body decay processes. Our analysis yields non-zero $A_{CP}$ in Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis and Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis. Considering its error, the observable CP violation can be observed at the order of $O(10^{-2})$. We strongly advocate for measuring these processes.
Note that the most ideal condition for measuring CPV is to produce charmed baryon and anti-baryon pairs simultaneously. Both the BESIII and Belle II Collaborations meet these conditions due to the most data on $\Lambda_{c}\bar{\Lambda}_{c}$ and $\Xi_{c}\bar{\Xi}_{c}$ pairs. For $\Lambda_{c}^{+}\to p\eta/n\pi^{+}$, their branching ratio has been measured by the BESIII and Belle II Collaborations. Therefore, it hold the potential for observing CP violation in charmed baryon decays at these two experimental facilities for the first time.
Acknowledgements- We thank Prof. Xiao-Gang He and Prof. Wei Wang for useful discussion. The work of Jin Sun is supported by IBS under the project code, IBS-R018-D1. The work of Zhi-Peng Xing is supported by NSFC under grant No.12375088 and No. 12335003. The work of Ruilin Zhu is supported by NSFC under grant No. 12322503 and No. 12075124, and by Natural Science Foundation of Jiangsu under Grant No. BK20211267.

### v3

@@ Abstract
CP violation in baryon decays has recently garnered significant attention. Traditionally, CP violation in charmed baryon decays is predicted to be at the order of $O(10^{-4})$. Utilizing recent measurements from BESIII, Belle and Belle II experiments, we conduct a comprehensive global analysis to estimate the contribution of topological diagrams in these processes. Our study reveals that topological diagrams alone cannot fully account for the experimental data. By incorporating new effects that are not included in traditional topological diagrams and exploring the possible source of the new effect, our analysis reveals the potential for significant CP violation in $\Lambda_{c}^{+}\to n\pi^{+}$ and $\Xi_{c}^{+}\to\Xi^{0}K^{+}$. This suggests a promising opportunity to observe CP violation for the first time in charmed baryon decays.
@@
In our analysis, we provide a global fit to determine all the IRA amplitudes in Ref. Xing:2023dni . With the help of the equivalence between IRA and TDA, we can not only test the correctness of the topological diagrams but also estimate the value of each diagram in TDA. As a result, we find that although the topological diagrams in TDA are basically consistent with our IRA numerical results, the topological diagram which only includes the $O_{1,2}$ operator cannot perfectly explain our IRA results. Without loss of generality, we can use a new diagram to express the conflicting parts of IRA and TDA and include them in our fit. To analyze the possible CPV introduced by the new diagram, we explored two possible source of the new diagram: SU(3) symmetry breaking and penguin operator and determine their contribution respectively. Our analysis shows that the new diagram may introduce a large CPV and it is promising to be observed in future experiments.
The global analysis of anti-triplet charmed baryon two-body decays with IRA- Using SU(3) flavor symmetry, the anti-triplet charmed baryon, light baryon octet, and light meson octet can be expressed by $3\times 3$ matrices $T_{c\bar{3}}$, $T_{8}$, and $P$, respectively Xing:2023dni . The SU(3) invariant decay amplitudes for anti-triplet charmed baryon two-body decays are constructed as $a_{15}\times\epsilon_{imn}(T_{c\bar{3}})^{[mn]}(H_{15})^{\{ik\}}_{j}(\overline{T_{8}})^{j}_{k}P^{l}_{l}$ and $a_{6}\times(T_{c\bar{3}})^{[ik]}(H_{\overline{6}})_{\{ij\}}(\overline{T_{8}})^{j}_{k}P^{l}_{l}$. Here the effective weak interaction Hamiltonian with IRA method is decomposed: $3\otimes\bar{3}\otimes 3=3\oplus 3\oplus\bar{6}\oplus 15$. The nonzero matrix are $H_{\bar{6}}$ and $H_{15}$, which are expressed in Ref. Xing:2023dni . By enumerating all possible matrix combinations, nine independent amplitudes can be constructed as $\{a,b_{6},c_{6},d_{6}\}$ and $\{a^{\prime},b_{15},c_{15},d_{15},e_{15}\}$. Using these nine SU(3) amplitudes, all amplitudes of anti-triplet charmed baryon two-body decays can be expressed.
@@
The assumptions are approximately supported by the numerical results within $1\sigma$ in Table. 1 and it shows the diagram $T_{1,2}$ actually give the main contribution. However, some form factors contradict the assumption, such as $|f_{1,15}|-|f_{2,16}|\sim 0.06$, $|f_{T_{17,18}}|-|f_{19}|\sim 0.02$ and $|g_{17,19}|-|g_{18}|\sim 0.05$. To focus on the largest effects, we ignore the discrepancy involving $f_{19}$ and primarily consider $|f_{1,15}|>|f_{2,16}|$ and $|g_{17,19}|>|g_{18}|$ in the subsequent analysis. This implies that there are other effects in the topology diagrams. For exploring the possible new effects, one defines the form factors $\textbf{f}_{b}$ and $\textbf{g}_{b}$ to absorb these effect as $\textbf{f}_{b}=f_{T_{1}}-f_{T_{2}}$ and $\textbf{g}_{b}=-g_{18}-g_{T_{7}}$. The corresponding topology diagrams $T_{8,9}$ are given in Fig. 1. Then the IRA form factors are expressed by TDA form factors and $\textbf{f}_{b},\textbf{g}_{b}$ as
@@
$g^{c}_{6}=g_{T_{6}},\quad g^{d}_{6}=3g_{T_{1}}-g_{T_{7}}+\frac{\textbf{g}_{b}}{2},\quad g^{d}_{15}=g_{T_{6}}.$ (5)
Since all the diagrams induced by $O_{1,2}$ are included in TDA diagrams $T_{1-7}$, it is reasonable to believe the diagrams $T_{8,9}$ are induced by SU(3) symmetry breaking or other operators like the penguin operators $O_{3-6}$. In Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis, we find that $g^{c}_{15}=0$, which is consistent with our global fit (Case I) in Table. 1. One can also directly observe that $f^{b}_{6}$ depends only on new effect form factors, which provides us an ideal condition to study the contribution from new effects.
@@
form factors Case II ($\chi^{2}$/d.o.f=0.875)
@@
vector(f) $f^{a}=0.0353(67)$ $f^{c}_{6}=-0.016(14)$ $f^{d}_{6}=-0.009(15)$ $f_{T_{1}}=0.009(14)$ $f_{T_{4}}=-0.005(16)$
$f^{a\prime}=-0.006(32)$ $\mathcal{R}e(\textbf{f}_{b}^{S})=-0.007(14)$ $\mathcal{R}e(\textbf{f}_{b}^{P})=0.00060(20)$ $\mathcal{I}m(\textbf{f}_{b}^{S})=0.003(16)$ $\mathcal{I}m(\textbf{f}_{b}^{P})=0.0013(18)$
$f^{b}_{15}=-0.021(12)$ $f^{c}_{15}=0.0047(52)$ $f^{d}_{15}=0.015(16)$ $f_{T_{6}}=0.020(21)$ $f_{T_{7}}=-0.012(33)$
$f^{e}_{15}=0.012(12)$
axial-vector(g) $g^{a}=0.003(29)$ $g^{b}_{6}=-0.035(18)$ $g^{c}_{6}=0.034(37)$ $g_{T_{1}}=0.019(34)$ $g_{T_{3}}=0.13(76)$
$g^{d}_{6}=-0.041(71)$ $g^{a\prime}=-0.06(50)$ $g_{T_{4}}=-0.104(84)$ $g_{T_{6}}=-0.026(51)$
$g^{b}_{15}=0.188(59)$ $g^{c}_{15}=0.022(41)$ $g^{d}_{15}=0.013(51)$ $g_{T_{7}}=-0.014(71)$
$g^{e}_{15}=0.006(53)$
@@
According to Eq.Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis, the IRA form factor $f^{b}_{6}$ depends solely on the new effect, potentially leads to significant CP violation. To study the CPV, the possible sources of the new form factor $\textbf{f}_{b}$ are discussed. Since the form factor arises from the difference between topological diagrams $T_{1}$ and $T_{2}$ in Fig.1, the most likely source is the SU(3) symmetry breaking effect. Especially when long-distance interactions enter the interaction vertex in Fig.1, the flavor symmetry breaking effect must be involved. Additionally, the penguin operator in Eq. 1 can also contribute to the new form factor $\textbf{f}_{b}$. When only flavor symmetry is considered, the difference in the Lorentz structure at the Hamiltonian vertex can be neglected. Consequently, the penguin operator has the same Hamiltonian matrix structure with the $O_{1,2}$ in TDA. However, in IRA, the penguin operator only contributes to $H_{3}$, which is neglected in our work. Therefore, the new form factor $\textbf{f}_{b}$ introduced by the conflicting parts of IRA and TDA must be considered.
According to the above analysis, the form factor can be divided into two parts $\textbf{f}_{b}=\textbf{f}^{S}_{b}+\textbf{f}^{P}_{b}e^{i\phi^{P}}$, where $\textbf{f}^{S}_{b}$ comes from the SU(3) symmetry breaking and $\textbf{f}^{P}_{b}e^{i\phi^{P}}$ represents the penguin operator contribution with its weak phase $\phi^{P}=-1.147\pm 0.026$. Including the two part of $\textbf{f}_{b}$, the form factors in our global analysis are defined as
$A^{q}_{6,15}=e^{i\phi_{1}}\bigg{(}\mathcal{R}e(A^{q}_{6,15})+\mathcal{I}m(A^{q}_{6,15})\bigg{)},$
$f^{b}_{6}=\frac{3}{2}\bigg{(}e^{i\phi_{1}}(\mathcal{R}e(\textbf{f}^{S}_{b})+\mathcal{I}m(\textbf{f}^{S}_{b}))$
$\quad\quad+e^{i\phi^{P}}(\mathcal{R}e(\textbf{f}^{P}_{b})+\mathcal{I}m(\textbf{f}^{P}_{b}))\bigg{)},$
$f^{d}_{6}=e^{i\phi_{1}}(\mathcal{R}e(f^{d}_{6})+\mathcal{I}m(f^{d}_{6}))$
$\quad\quad-e^{i\phi^{P}}\frac{3}{2}(\mathcal{R}e(\textbf{f}^{P}_{b})+\mathcal{I}m(\textbf{f}^{P}_{b})),$
$f^{e}_{15}=e^{i\phi_{1}}(\mathcal{R}e(f^{e}_{15})+\mathcal{I}m(f^{e}_{15}))$
$\quad\quad-e^{i\phi^{P}}\frac{3}{2}(\mathcal{R}e(\textbf{f}^{P}_{b})+\mathcal{I}m(\textbf{f}^{P}_{b})),\;A=f,g,$ (6)
where $q=a,b,c,d,e$. The weak phase $\phi_{1}$ arises from the current-current operator as $\phi_{1}=\arg(V_{cq}^{*}V_{uq^{\prime}})\approx 0,-\pi$. On can find that the penguin operator only contribute to the Cabibbo-suppressed processes. The SU(3) symmetry requires the form factor corresponding to different processes to be equal Geng:2023pkr ; Zhong:2024qqs , one has the chance to determine the form factor $\textbf{f}^{S}_{b}$ by the Cabibbo-allowed and doubly Cabibbo-suppressed processes. Based on the symmetry breaking form factor we derived, the penguin contribution form factor $\textbf{f}^{P}_{b}$ can be further determined by the Cabibbo-suppressed processes.
@@
By introducing the weak phase of penguin operator $\phi^{T}$, one can derive the CPV if the global fit gives a nonzero value of $\textbf{f}^{P}_{b}$. In our analysis, we obtains the fit results with $\chi^{2}/d.o.f=0.875$, indicating that our fit is reasonable and the form factors effectively explain the experimental data.
@@
With the determined penguin operator contribution $\mathcal{R}e(\textbf{f}^{P}_{b})=0.0006(2)$ and $\mathcal{I}m(\textbf{f}^{P}_{b})=0.0013(18)$, one can predict the CP violation ($A_{CP}$) for the Cabibbo-suppressed processes. Considering the ratio of the form factor from current-current and penguin operator $\frac{\langle O_{1,2}\rangle}{\langle O_{3-6}\rangle}\sim O(10^{-2})$, our study implies that the contribution of the QCD penguin may be more substantial than conventionally assumed, necessitating further theoretical investigation in the future.
@@
Based on the fitted results, we predict the values $A_{CP}$ for the Cabibbo-suppressed processes and found that almost all the predicted $A_{CP}$ values are approximately zero within errors, except for the following processes
$A_{CP}^{\Lambda_{c}^{+}\to n\pi^{+}}=0.076(57),\;A_{CP}^{\Xi_{c}^{+}\to\Xi^{0}K^{+}}=0.092(70).$ (7)
We strongly recommend the experimental measurement of the processes: $\Lambda_{c}^{+}\to n\pi^{+}$ since their branching ratio has been measured.
@@
By leveraging the equivalence between IRA and TDA methods, our global analysis extends to complex form factors. By considering the SU(3) symmetry breaking and penguin operator as the source of form factor $\textbf{f}_{b}$, the corresponding contributions of SU(3) symmetry breaking and penguin operator are determined through global analysis respectively. Then we further explore CP violation in anti-triplet charmed baryon two-body decay processes. Our analysis yields non-zero $A_{CP}$ in Eq. 7. Considering its error, the observable CP violation can be observed at the order of $O(10^{-2})$. We strongly advocate for measuring these processes.
Note that the most ideal condition for measuring CPV is to produce charmed baryon and anti-baryon pairs simultaneously. The BESIII Collaborations meet these conditions due to the most data on $\Lambda_{c}\bar{\Lambda}_{c}$ and $\Xi_{c}\bar{\Xi}_{c}$ pairs. For $\Lambda_{c}^{+}\to n\pi^{+}$, their branching ratio has been measured by the BESIII Collaborations. Therefore, it hold the potential for observing CP violation in charmed baryon decays at this experimental facilities for the first time.
Acknowledgements.
We thank Prof. Xiao-Gang He, Prof. Wei Wang and Prof. YuJi Shi for useful discussion. The work of Jin Sun is supported by IBS under the project code, IBS-R018-D1. The work of Ruilin Zhu is supported by NSFC under grant No. 12322503 and No. 12075124, and by Natural Science Foundation of Jiangsu under Grant No. BK20211267. The work of Zhi-Peng Xing is supported by NSFC under grant No.12375088 and No. 12335003.